OpenStudy (anonymous):

x^-8/y^-7 * y^-7/x^-8 evaluate

OpenStudy (anonymous):

what happens when you multiply reciprocals?

OpenStudy (anonymous):

hint.. \[\huge \frac{x^a}{x^b}=x^{a-b}\]

OpenStudy (anonymous):

idk thats y im asking can you solve it for me

OpenStudy (anonymous):

what have you tried?

OpenStudy (lgbasallote):

Hint: 2 x 1 = 1 - 2

OpenStudy (anonymous):

idk how to do it period

OpenStudy (lgbasallote):

you have to try..we're not homework answerers @9013955 :) we are here to teach. please do understand ^_^ so plase listen to the hints

OpenStudy (anonymous):

just try... it doesn't matter if it's wrong ... then show us what have you tried ... and we'll help you ! ... How can we help you when we don't even know where are you stuck !